OpenStudy (anonymous):

1 2/3ft & 2/3ft what is the area of this rectangle ?

OpenStudy (texaschic101):

area of a rectangle is width x height. so we basically have 1 2/3 x 2/3 turn the 1 2/3 into an improper fraction.. 1 2/3 = (1 x 3 + 2)/3 = 5/3 now multiply... 5/3 x 2/3 = 10/9 any questions ?

OpenStudy (texaschic101):

or if you want it in mixed number form... 10/9 = 1 1/9

OpenStudy (anonymous):

which one would it be then ? here are the answers ; 160in sq2 , 56in sq2 , 28in sq2 , 80in sq2 or none of the above ?